How much energy in KJ is needed to raise the temperature of a 55g sample with a specific heat 0.897 J/gk from 22.4* C to 94*C?


1
Expert's answer
2022-04-05T02:27:03-0400

∆H= MC∆T

M= 55g

C= 0.897J/g/K

∆T=94-22.4= 71.6K

∆H= 55×0.897×71.6

= 3532.386J

=3532.386÷1000 kJ

= 3.532386kJ

= 3.532kJ